OpenStudy (anonymous):

Which of the following is the slope between the two points (0, -4) and (-2, 5)? 9 - --- 2 2 -- 9 2 - --- 9 9 - 2

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

Slope of two points = (y2-y1) / (x2-x1) (x1,y1) = (0,-4) and (x2,y2) = (-2,5) So, slope = 5-(-4) / -2-0 = 5+4 / -2 = 9/-2 = -9/2

OpenStudy (anonymous):

Its A!!

OpenStudy (anonymous):

oh, my bad thanks

OpenStudy (anonymous):

: )